Pb(NO3)2 + 2 NaI à PbI2 + 2 NaNO                                 

1)     How many moles of lead (II) nitrate are needed to react completely with 10 moles of sodium iodide?

1
Expert's answer
2022-03-04T05:12:03-0500

Balanced equation is;

Pb(NO3)2 + 2 NaI = PbI2 + 2 NaNO3

From the balanced equation, for every 2 moles of sodium iodide, there is 1 mole of lead (ii) nitrate.

So, the moles needed to react completely with 10 moles of sodium iodide is; (10÷2) = 5

= 5 moles of lead (ii) nitrate